Accelerator Update for September 1, 2008 - September 3, 2008
Monday Sept 1
The day shift began with Operations monitoring the Tevatron (TeV) store 6391 (102.24E30) and the Recycler (RR) stash (230.7E10), with the Antiproton Source (Pbar) stacking (22.7E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the MiniBooNE and NuMI experiments taking beam.
At 8:55 AM, the P2 transfer line beam permit tripped off. It reset.
Pbar resumed stacking at 8:59 AM. A Pbar expert reported that he had been running a program that had sent a bad setting for the kicker’s timing.
At 12:17 PM, a Linac Klystron Debuncher cavity alarmed on a flow fault for the first of three times in 28 minutes. It reset.
At 1:32 PM, a Linac Klystron Debuncher cavity alarmed. It reset.
Operations terminated store 6391 at 1:39 PM.
At 1:59 PM, a Linac Klystron Debuncher cavity alarmed. It reset.
Operations put the TeV through a dry squeeze at 4:03 PM.
Operations reported at 4:03 PM that the Recycler had stashed antiprotons five times during the day shift.
Operations put the TeV into shot setup at 4:40 PM.
Operations established store 6393 at 6:40 PM with an initial luminosity of 252.39E30.
At 7:19 PM, the Linac 400MeV area beam permit tripped off due to bad beam positions. It reset.
At 7:56 PM, the H- chopper power supply tripped off for the first of two times. Operators investigated and found a power supply had failed. Operations called in an expert to replace the power supply.
At 9:10 PM, BRF10 tripped off. It reset.
Linac resumed sending beam at 9:56 PM.
At 11:15 PM, Operations reported at vacuum burst in the MI-8 transfer line. Everything reset.
Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.
Tuesday Sept 2
The midnight shift began with Operations monitoring store 6393 (124.77E30) and the stash (132.3E10), with the Pbar stacking (33.1E10), and with the MiniBooNE and NuMI experiments taking beam.
At 5:57 AM, the TeV suffered another quench, this time at sector A4. Later, experts reported that this quench was due to a lead flow problem.
At 6:11 AM, a Booster vacuum pump (IPRF23) tripped off and caused BRF14 and 15 to also trip. Everything reset.
Operations reported at 8:13 AM that the Recycler had stashed five times during the midnight shift.
An RF expert brought MIRF5 back online at 8:20 AM.
Operations put the TeV through a QBS ramp test at 9:35 AM.
TeV experts began a study at 11:27 AM.
The Neutron Therapy Facility reported at 12:07 PM that they had treated their last patient for the day.
TeV studies ended at 1:18 PM.
The NuMI beam permit tripped off due to a MI RF problem at 1:24 PM.
At 1:57 PM, a Linac quadrupole power supply (QPS225) breaker tripped off for the first of two times in 10 minutes. It reset.
Operations put the TeV into shot setup at 2:28 PM.
At 2:37 PM, LRF1 tripped off. It reset.
At 2:38 PM, BRF9 tripped off. It reset.
At 2:47 PM, Operations reported that the Booster West Anode power supply had tripped off with no indication as to why, as did an Accumulator RF station (ARF1). Everything reset.
At 2:50 PM, QPS225 tripped off for the first of two times in 13 minutes. Operators investigated. Linac resumed sending beam at 3:05 PM.
At 3:14 PM, the Linac Klystron Cavity water skid tripped off. It reset.
At 3:40 PM, Operations reported a vacuum pump problem for the Linac RF buncher. Operations contacted experts.
Operations reported at 3:59 PM that the Recycler had stashed three times during the day shift.
Operations reported at 4:33 PM that MI RF experts were changing the modulator for MIRF5.
Operations established store 6397 at 4:42 PM with an initial luminosity of 300.33E30.
A MI expert reported at 4:46 PM that he was replacing a filtered voltage chassis for the MI-40 lower power supplies.
RF experts reported at 5:05 PM that they had completed their work on MIRF5.
MI resumed sending beam at 5:14 PM.
At 6:43 PM, Operations reported a couple of NuMI trips due to MIRF2. A RF expert investigated.
At 10:18 PM, LRF4 tripped off. It reset.
At 11:07 PM, Operations reported a couple more NuMI trips, this time due to MIRF6.
Operations reported at midnight that the Recycler had stashed five times during the evening shift.
Wednesday Sept 3
The midnight shift began with Operations monitoring store 6397 (120.34E10) and the stash (185.0E10), with the Pbar stacking (35.5E10), and with the MiniBooNE and NuMI experiments taking beam.
At 1:41 AM, Operations reported that MIRF6 had caused several NuMI trips. After 30 more minutes of trips, operators bypassed the station.
At 4:57 AM, KRF6 tripped off on a LCW flow fault. It reset.
At 5:30 AM, the Recycler Pelletron tripped off. It took several tries to reset, but eventually it came back online.
At 5:34 AM, the lithium lens tripped off for no obvious reason. It reset.
At 7:33 AM, the Klystron Debuncher tripped on a cavity flow fault. The flow meter would not reset. Operations contacted a water expert and a Linac expert.
Operations put the Pbar into standby at 7:43 AM.
Operations reported at 8:02 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 09/03/2008 email Fermilab|
|Security, Privacy, Legal|